OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative sixth power

OpenStudy (anonymous):

\[\large \frac{1}{\sqrt[3]{x^{-6}}}\]?

OpenStudy (anonymous):

yes that's what it looks like

OpenStudy (anonymous):

the negative exponent means flip it, so you get \[\sqrt[3]{x^6}\] as a first step

OpenStudy (anonymous):

as a rational exponent this is \[x^{\frac{6}{3}}\] and since \(\frac{6}{3}=2\) your final answer is \(x^2\)

OpenStudy (anonymous):

thank you !! :)

OpenStudy (anonymous):

yw